OpenStudy (anonymous):

Integrate. cos^2x tan^3x

OpenStudy (lgbasallote):

hint: \[\tan^3 x \implies \frac{\sin^3 x}{\cos^3 x}\]

OpenStudy (anonymous):

Yeah I think I got it.

OpenStudy (lgbasallote):

congrats :D

OpenStudy (anonymous):

I got up to \[\int\limits_{}^{}(1/2-1/2\cos2x)\sin(x) \cos ^{-1} (x)\]

OpenStudy (anonymous):

Yeah, now I'm lost.

OpenStudy (anonymous):

oh got it's \[\int\limits_{}^{}(1-\cos^2 x)\sin(x) \cos ^{-1} x\] u=cosx

OpenStudy (zarkon):

\[\cos^2(x)\frac{\sin^3 x}{\cos^3 x}=\frac{\sin^3(x)}{\cos(x)}=\sin(x)\frac{\sin^2(x)}{\cos(x)}\] \[=\sin(x)\frac{1-\cos^2(x)}{\cos(x)}\] ...yep

OpenStudy (anonymous):

Thanks!